OpenStudy (anonymous):

im stuck (x^2y/xy^3)^2

OpenStudy (imstuck):

can you write that in the equation editor or using the drawing tool? All those ^'s and parenthesis are really confusing as to where they actually go and what they mean.

OpenStudy (dangerousjesse):

I'm guessing they meant \(\frac{x^2 y}{xy^3}^2\)

OpenStudy (anonymous):

\[(\frac{ x ^{2}y }{ xy^{3} })^{2}\] I think he meant that?

OpenStudy (dangerousjesse):

For all exponents, \(a^n a^m = a^{n+m}\). Apply this to \(\frac{x^2 y}{x y^3}\). Combine powers. \(\frac{x^2 y}{x y^3} = x^{2-1} y^{1-3}\): \((x^2-1 y^1-3)^2\) Evaluate 2-1. \(2-1 = 1:\) \((x y^{1-3})^2\) Evaluate \(1-3\). \(1-3 = -2:\) \((x y^{-2})^2\) Distribute exponents over quotients in \((\frac{x}{y^2})^2.\) Multiply each exponent in \(\frac{x}{y^2}\) by 2: \(x^2 y^{-2×2}\) Multiply 2 and -2 together. \(2 (-2) = -4:\)

OpenStudy (dangerousjesse):

That's what I meant, typo :P forgot the parentheses too haha